a+(a-3)=(a+4)-(a+1)= can you please show me step by step, I can't seem to figure out what I am doing wrong. Thank you
Answer by rfer(16322)

a+a-3=a+4-a-1
2a=4-1+3
2a=6
a=6/2
a=3